𝑇: ℝ3→ℝ3 defined by 𝑇(π‘₯1,π‘₯2,π‘₯3)=(π‘₯1+π‘₯2,π‘₯2+π‘₯3,π‘₯3+π‘₯1).


1
Expert's answer
2021-02-01T01:38:05-0500

"T=\begin{pmatrix}
 1 & 1 & 0\\
 0 & 1 & 1\\
 1 & 0 & 1
\end{pmatrix}\\
det(T)=1\begin{vmatrix}
 1 & 0 \\
 0 & 1
\end{vmatrix}-1\begin{vmatrix}
 0 & 1 \\
 1 & 1
\end{vmatrix}+0\\
det(T)=2,\\
 \text{then, T is invertible}\\
\text{Augmented matrix of T is,}\\
\begin{pmatrix}
 1 & 1 & 0 & & 1 & 0 & 0\\
 0 & 1 & 1 & & 0 & 1 & 0\\
 1 & 0 & 1 & & 0 & 0 & 1
\end{pmatrix} \\-R_1+R_3 \implies R_3\\
\begin{pmatrix}
 1 & 1 & 0 & & 1 & 0 & 0\\
 0 & 1 & 1 & & 0 & 1 & 0\\
 0 & -1 & 1 & & -1 & 0 & 1
\end{pmatrix} \\R_3+R_1 \implies R_1\\
\begin{pmatrix}
 1 & 0 & 1 & & 0 & 0 & 1\\
 0 & 1 & 1 & & 0 & 1 & 0\\
 0 & -1 & 1 & & -1 & 0 & 1
\end{pmatrix} \\R_2+R_3 \implies R_3\\
\begin{pmatrix}
 1 & 0 & 1 & & 0 & 0 & 1\\
 0 & 1 & 1 & & 0 & 1 & 0\\
 0 & 0 & 2 & & -1 & 1 & 1
\end{pmatrix} \\ \frac{1}{2}R_3\\
\begin{pmatrix}
 1 & 0 & 1 & & 0 & 0 & 1\\
 0 & 1 & 1 & & 0 & 1 & 0\\
 0 & 0 & 1 & & -\frac{1}{2} & \frac{1}{2} & \frac{1}{2} 
\end{pmatrix} \\
-R_3+R_1 \implies R_1\\
-R_3+R_2 \implies R_2\\
\begin{pmatrix}
 1 & 0 & 0 & & \frac{1}{2} & -\frac{1}{2} & \frac{3}{2} \\
 0 & 1 & 0 & & \frac{1}{2} & \frac{3}{2} & -\frac{1}{2}\\
 0 & 0 & 1 & & -\frac{1}{2} & \frac{1}{2} & \frac{1}{2} 
\end{pmatrix} \\

T^{-1}=\begin{pmatrix}
\frac{1}{2} & -\frac{1}{2} & \frac{3}{2} \\
\frac{1}{2} & \frac{3}{2} & -\frac{1}{2}\\
 -\frac{1}{2} & \frac{1}{2} & \frac{1}{2} 
\end{pmatrix} \\"